# Resilient by Design – Active Inference for Distributed Continuum Intelligence

### v1

@@ IV Theoretical Results and Discussions
This section presents the theoretical foundations demonstrating that the proposed PAIR-Agent framework achieves resilient operation in the DCC through probabilistic reasoning, causal inference, and active self-healing. Three main results establish its correctness and convergence properties.
Result 1 (Convergence of Belief Updating). Let $P^{t}(f_{i})$ denote the prior fault belief at checkpoint $t$ and $Q^{t+1}(f_{i})$ its posterior after processing new evidence $\Delta\mathcal{D}^{t+1}$. Under the Bayesian update rule $Q^{t+1}(f_{i})\propto P(\Delta\mathcal{D}^{t+1}\mid f_{i})P^{t}(f_{i}),$ the Kullback–Leibler divergence between successive posterior estimates satisfies $D_{\mathrm{KL}}\!\left(Q^{t+1}(f_{i})\parallel Q^{t}(f_{i})\right)\to 0\quad\text{as }|\Delta\mathcal{D}^{t+1}|\to 0.$ Hence, as log updates become incremental, the posterior distribution converges almost surely to the true fault probability $P^{*}(f_{i})$. This ensures consistent belief refinement across checkpoints, guaranteeing stable and bounded fault estimation in the presence of streaming data.
Result 2 (Optimal Action Selection under Expected Free Energy Minimization). Let $G(a_{t})$ denote the expected free energy associated with action $a_{t}\in\mathcal{A}$: $G(a_{t})=\mathbb{E}_{Q(s_{t+1}\mid a_{t})}\Big[D_{\mathrm{KL}}\!\left(Q(s_{t+1})\parallel P(s_{t+1})\right)-\ln P(o_{t+1}\mid s_{t+1})\Big].$ Assuming $G(a_{t})$ is convex in the neighborhood of the optimal policy and both $Q(s_{t+1}\mid a_{t})$ and $P(s_{t+1})$ are Lipschitz continuous, the selected action $a_{t}^{*}=\arg\min_{a_{t}}G(a_{t})$ minimizes the upper bound of predictive uncertainty while maximizing the evidence of expected observations. Thus, PAIR-Agent’s policy converges to an equilibrium minimizing both epistemic (uncertainty-driven) and pragmatic (goal-driven) free energy, ensuring that corrective actions asymptotically restore the DCC to its desired operational manifold.
Result 3 (System-Level Resilience via Active Healing Loop). Consider the system state dynamics $s_{t+1}=f(s_{t},a_{t},\xi_{t})$, where $\xi_{t}$ represents random disturbances or failures. Define resilience as the probability of maintaining bounded deviation from the nominal state, i.e., $\mathbb{P}(\|s_{t+1}-s^{*}\|\leq\epsilon)\geq 1-\delta,$ for tolerance $\epsilon$ and confidence $1-\delta$. Under PAIR-Agent’s closed-loop policy that minimizes both expected free energy $G(a_{t})$ and residual fault probability $\sum_{i}\hat{P}(f_{i}\mid a_{t})$, the composite Lyapunov function $V_{t}=F^{t}+\lambda\sum_{i}\hat{P}(f_{i})$ is non-increasing for all $t>0$, the mean-square boundedness of $s_{t}$. Therefore, the DCC remains resilient against stochastic faults, achieving self-stabilization through the proposed PAIR-Agent.
@@ V Conclusions
This paper introduced the PAIR-Agent as a unified framework for achieving resilience in DCC systems. The PAIR-Agent agent performs: (i) constructing a CFG through device log analysis, (ii) identifying faults while managing certainties and uncertainties using Markov blankets and the free-energy principle, and (iii) autonomously healing issues via active inference. Through its closed-loop cycle of observation, inference, and action, PAIR-Agent ensures adaptive stability, self-healing capability, and sustained operational continuity, marking a significant step toward fully autonomous and resilient DCC systems. Theoretical results confirm its reliability, and future work will extend this research with detailed and distributed implementation, empirical validation, and evaluation under real-time environments.

### v2

@@ IV Theoretical Results and Discussions
This section presents the theoretical foundations showing how PAIR-Agent achieves resilient operation in the DCC through principled probabilistic inference and cautious, uncertainty-aware action selection. The results highlight four key properties: the locality and scalability of fault inference, the quality of the variational posterior approximation, the safety of the agent’s corrective actions, and the robustness of belief updates under missing or delayed logs.
Result 1 (Locality and Scalability of Fault Inference). For any fault variable $f_{i}$, PAIR-Agent only needs the variables in its Markov blanket $\mathcal{M}(f_{i})$ (its parents, children, and co-parents) to update the belief $Q(f_{i})$. Thus, the time and memory per fault update depend on $|\mathcal{M}(f_{i})|$, and not on the total number of variables in the CFG. If the average blanket size is $B$, updating all $d$ faults in one round costs $O(d\cdot B)$, which stays manageable as the DCC grows, provided local neighborhoods stay small. This guarantees that the pipeline in Fig. 1 scales due to inference being local by design.
Result 2 (Best Possible Approximate Posterior). At checkpoint $t+1$, PAIR-Agent selects $Q^{t+1}(\mathbf{f})$ by minimizing the variational free energy $F^{t+1}(Q)$. Since, $F^{t+1}(Q)$ differs from $D_{KL}(Q(\mathbf{f})\parallel P^{t+1}(\mathbf{f}\mid\mathbf{x}^{t+1}))$ only by a constant independent of $Q$ (i.e., $\ln P^{t+1}(\mathbf{x}^{t+1})$), the update $Q^{t+1}=\operatorname*{\arg\min}_{Q\in\mathcal{Q}}F^{t+1}(Q)$ is equivalently the closest distribution to the exact Bayesian posterior within the chosen variational family $\mathcal{Q}$. Therefore, given the current CFG and log data, PAIR-Agent computes the best posterior approximation it can represent, ensuring principled and internally consistent fault inference.
Result 3 (Conservative Action Selection [Never Worse Than Doing Nothing]). Assume that the action set includes a baseline ”do nothing” action $a^{\varnothing}$, which keeps the current system configuration. Then, because PAIR-Agent selects $a_{t}^{*}=\operatorname*{\arg\min}_{a_{t}}G(a_{t})$, the chosen action never has higher expected free energy than $a^{\varnothing}$. Conversely, when an action is predicted to move $Q(\mathbf{x}\mid a)$ closer to preferences and/or reduce uncertainty, it will be selected. This guarantees that the agent will not take an action that it predicts will diverge the system from its operational objectives and not increase its model certainty.
Result 4 (Robustness to Missing or Delayed Logs). At checkpoint $t+1$, the belief update for any fault $f_{i}$ depends only on variables in its Markov blanket $\mathcal{M}(f_{i})$. Therefore, missing or delayed logs for variables outside $\mathcal{M}(f_{i})$ do not change the updated belief $Q^{t+1}(f_{i})$. Since telemetry in DCC is often partial or out-of-order, this property makes PAIR-Agent’s inference insensitive to unrelated data gaps and supports asynchronous, scalable updates: only logs inside the small local neighborhood $\mathcal{M}(f_{i})$ can affect the diagnosis of $f_{i}$.
@@ V Conclusions
This paper introduced the PAIR-Agent as a unified framework for achieving resilience in DCC systems. The PAIR-Agent agent performs: (i) constructing a CFG through device log analysis, (ii) identifying faults while managing uncertainties using Markov blankets and the FEP, and (iii) autonomously healing issues via active inference. Through its closed-loop cycle of observation, inference, and action, PAIR-Agent ensures adaptive stability, self-healing capability, and sustained operational continuity, marking a significant step toward autonomous and resilient DCC systems. Theoretical results confirm its scalability, principled and accurate fault inference, safe decision-making, and robustness to incomplete telemetry. Future work will extend this research with detailed and distributed implementation, empirical validation, and evaluation under a real-time testbed.